# Linear Complementary Pairs of Quasi-Cyclic and Quasi-Twisted Codes

### v1

@@ Conclusion
Linear complementary pairs (LCP) of codes generalize linear complementary dual (LCD) codes and have applications in resistance against side-channel and fault-injection attacks. In this work, we presented a polynomial characterization of LCPs of quasi-cyclic and quasi-twisted codes of index $2$. As a consequence, we provided necessary and sufficient conditions for LCPs of one-generator quasi-cyclic and quasi-twisted codes. In addition, we provided several examples of LCPs of quasi-cyclic and quasi-twisted codes with (almost) optimal security parameters.
@@ Acknowledgment
K. Abdukhalikov was supported by UAEU grants G00004233 and G00004614. D. Ho was supported by the Tromsø Research Foundation (project “Pure Mathematics in Norway”) and UiT Aurora project MASCOT. G. K. Verma was supported by UAEU grant G00004614. S. Ling is supported by Nanyang Technological University Research Grant No. 04INS000047C230GRT01 and UAEU grant G00004233.

### v2

@@ Conclusion
Linear complementary pairs (LCP) of codes generalize linear complementary dual (LCD) codes and have applications in resistance against side-channel and fault-injection attacks. In this work, we presented a polynomial characterization of LCPs of quasi-cyclic and quasi-twisted codes of index $2$. As a consequence, we provided necessary and sufficient conditions for LCPs of one-generator quasi-cyclic and quasi-twisted codes. In addition, we provided several examples of LCPs of quasi-cyclic and quasi-twisted codes with optimal security parameters.
@@ Acknowledgment
K. Abdukhalikov and G. K. Verma were supported by UAEU grants G00004233 and G00004614. D. Ho was supported by the Tromsø Research Foundation (project “Pure Mathematics in Norway”), and UiT Aurora project MASCOT. S. Ling is supported by Nanyang Technological University Research Grant No. 04INS000047C230GRT01.